News & Status
  • News & status not loaded yet.

Javascript is turned off in your web browser.

To take full advantage of this site, please turn it on and then refresh the page.

2017-01-10 (Zeiss LSM 700 Confocal): Zeiss LSM700 moving down the hall to Biophysics B3

The Microscope Facility Zeiss LSM700 will be relocated to Biophysics B3, just down the hall from its current location (Physiology G09) to make way for the new Zeiss GaAsP confocal. We anticipate that the 700 system will be offline from 10am-2pm Thursday (1/12) and again on Friday morning (1/13).

All certified 700 users will have access to the new room which has 24/7 swipe access. Please remember to have your badge when you come to image on the 700.

More information about transitioning users to the new microscope system, Zeiss GaAsP, will be posted soon. Thank you for your cooperation.

Microscope Facility News & Equipment Status

To change News view:
Two months
Quarter
Year
All
Select multiple with control-key
Restrict listing
to selected
equipment
(Exclude announcements)

2017-11-13 (Philips CM120 TEM): Philips TEM: chiller issues

The Philips TEM is offline. The High tension is not coming on and the ODP on the water chiller went down. The FEI engineer is scheduled for a visit on Tuesday. We apologize for the inconvenience.

2017-10-26 (Hitachi 7600 TEM): Hitachi TEM, back online, chiller stabilized

10/26/2017 12:26pm (UPDATE)>> Back online, chiller stabilized
The Hitachi water chiller has been serviced and is the Hitachi TEM is back online. Thank you for your patience.

****************************************
10/25/2017 4:21pm Hitachi TEM, chiller is down, scope unusable
The Hitachi chiller is down water is running too hot (114 F), I shut down the scope removed my sample (the sample holder is secured in my office). I went downstairs to ask for maintenance to look at it. They will contact Leroy in the morning. Scope is unusable. -Michael Delannoy

2017-10-13 (Zeiss LSM 510-meta #2): Zeiss 510-meta, computer/network issue, scope inoperable

The 510 computer operation system is corrupted and unrepairable by the Zeiss engineer. We are looking into other options (using a spare computer) and other attempts (BSNO service). Both are in action, but neither will be completed before next Monday.

2017-10-13 (Olympus Station 1 (inverted)): Olympus 1 (inverted): controller box not working, in repair until Mon Oct 16

The Olympus inverted microscope system is unavailable until Mon, Oct 16. It appears that liquid, which was spilled on the controller, leaked down into the electronic circuit board. This may have been causing a malfunction in the operation of the microscope. Scot has cleaned off the circuit board and will reinstall it when it is completely dry. I am sorry for last minute notice and for any inconvenience to you all.

Last modified